OpenStudy (anonymous):

y varies directly as the square of x. when x=4, y=8. find x when y=32

OpenStudy (anonymous):

y=kx^2, where k is a constant when x=4,y=8 8=k4^2 8=16k k=8/16=1/2 \[y=\frac{ 1 }{ 2 }x ^{2},when y=32,\] \[x ^{2}=2y=2*32=64,x=\pm8\]